OpenStudy (anonymous):

Complete the function table below for the function F(x) = 4x - 6. x 1 2 3 4 F(x) -2 The numbers in the second row of the completed table make an arithmetic sequence. A. True B. False @aum

OpenStudy (anonymous):

would it be false @aum

OpenStudy (aum):

F(x) = 4x - 6 when x = 1, F(x) = 4 * 1 - 6 = -2 when x = 2, F(x) = 4 * 2 - 6 = 2 when x = 3, F(x) = 4 * 3 - 6 = 6 when x = 4, F(x) = 4 * 4 - 6 = 10 The second row is: -2, 2, 6, 10 Th differences are: 2 - (-2) = 4; 6 - 2 = 4; 10 - 6 = 4 The differences are the same and equal to 4. Therefore, the second row is in arithmetic progression. So it is True.